libffi: a foreign function interface library

For Version 3.4.2 of libffi

Anthony Green

This manual is for libffi, a portable foreign function interface library. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2008–2019, 2021 Anthony Green and Red Hat, Inc. Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the “Software”), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions: The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software. THE SOFTWARE IS PROVIDED “AS IS”, WITHOUT WARRANTY OF ANY KIND, EXPRESS OR IMPLIED, INCLUDING BUT NOT LIMITED TO THE WARRANTIES OF MERCHANTABILITY, FITNESS FOR A PARTICULAR PURPOSE AND NONIN-FRINGEMENT. IN NO EVENT SHALL THE AUTHORS OR COPYRIGHT HOLDERS BE LIABLE FOR ANY CLAIM, DAMAGES OR OTHER LIABILITY, WHETHER IN AN ACTION OF CONTRACT, TORT OR OTHERWISE, ARISING FROM, OUT OF OR IN CONNECTION WITH THE SOFTWARE OR THE USE OR OTHER DEALINGS IN THE SOFTWARE.

1 What is libffi?

Compilers for high level languages generate code that follow certain conventions. These conventions are necessary, in part, for separate compilation to work. One such convention is the calling convention. The calling convention is a set of assumptions made by the compiler about where function arguments will be found on entry to a function. A calling convention also specifies where the return value for a function is found. The calling convention is also sometimes called the ABI or Application Binary Interface.

Some programs may not know at the time of compilation what arguments are to be passed to a function. For instance, an interpreter may be told at run-time about the number and types of arguments used to call a given function. ‘Libffi’ can be used in such programs to provide a bridge from the interpreter program to compiled code.

The ‘libffi’ library provides a portable, high level programming interface to various calling conventions. This allows a programmer to call any function specified by a call interface description at run time.

FFI stands for Foreign Function Interface. A foreign function interface is the popular name for the interface that allows code written in one language to call code written in another language. The ‘libffi’ library really only provides the lowest, machine dependent layer of a fully featured foreign function interface. A layer must exist above ‘libffi’ that handles type conversions for values passed between the two languages.

2 Using libffi

2.1 The Basics

‘Libffi’ assumes that you have a pointer to the function you wish to call and that you know the number and types of arguments to pass it, as well as the return type of the function.

The first thing you must do is create an ffi_cif object that matches the signature of the function you wish to call. This is a separate step because it is common to make multiple calls using a single ffi_cif. The cif in ffi_cif stands for Call InterFace. To prepare a call interface object, use the function ffi_prep_cif.

[Function] ffi_status ffi prep cif (ffi cif *cif, ffi abi abi, unsigned int nargs,

ffi type *rtype, ffi type **argtypes) This initializes cif according to the given parameters. abi is the ABI to use; normally FFI_DEFAULT_ABI is what you want. Section 2.4 [Multiple ABIs], page 11 for more information. nargs is the number of arguments that this function accepts. rtype is a pointer to an ffi_type structure that describes the return type of the function. See Section 2.3 [Types], page 3. argtypes is a vector of ffi_type pointers. argtypes must have nargs elements. If nargs is 0, this argument is ignored. ffi_prep_cif returns a libffi status code, of type ffi_status. This will be either FFI_OK if everything worked properly; FFI_BAD_TYPEDEF if one of the ffi_type objects is incorrect; or FFI_BAD_ABI if the abi parameter is invalid.

If the function being called is variadic (varargs) then ffi_prep_cif_var must be used

instead of ffi_prep_cif.

[Function] ffi_status ffi prep cif var (ffi cif *cif, ffi abi abi, unsigned int

   nfixedargs, unsigned int ntotalargs, ffi type *rtype, ffi type
   **argtypes)

This initializes cif according to the given parameters for a call to a variadic function. In general its operation is the same as for ffi_prep_cif except that: nfixedargs is the number of fixed arguments, prior to any variadic arguments. It must be greater than zero. ntotalargs the total number of arguments, including variadic and fixed arguments. argtypes must have this many elements. ffi_prep_cif_var will return FFI_BAD_ARGTYPE if any of the variable argument types are ffi_type_float (promote to ffi_type_double first), or any integer type small than an int (promote to an int-sized type first). Note that, different cif’s must be prepped for calls to the same function when different numbers of arguments are passed. Also note that a call to ffi_prep_cif_var with nfixedargs=nototalargs is NOT equivalent to a call to ffi_prep_cif.

Note that the resulting ffi_cif holds pointers to all the ffi_type objects that were used during initialization. You must ensure that these type objects have a lifetime at least as long as that of the ffi_cif.

To call a function using an initialized ffi_cif, use the ffi_call function:

[Function] void ffi call (ffi cif *cif, void *fn, void *rvalue, void **avalues)

This calls the function fn according to the description given in cif. cif must have already been prepared using ffi_prep_cif. rvalue is a pointer to a chunk of memory that will hold the result of the function call. This must be large enough to hold the result, no smaller than the system register size (generally 32 or 64 bits), and must be suitably aligned; it is the caller’s responsibility to ensure this. If cif declares that the function returns void (using ffi_type_void), then rvalue is ignored. In most situations, ‘libffi’ will handle promotion according to the ABI. However, for historical reasons, there is a special case with return values that must be handled by your code. In particular, for integral (not struct) types that are narrower than the system register size, the return value will be widened by ‘libffi’. ‘libffi’ provides a type, ffi_arg, that can be used as the return type. For example, if the CIF was defined with a return type of char, ‘libffi’ will try to store a full ffi_arg into the return value. avalues is a vector of void * pointers that point to the memory locations holding the argument values for a call. If cif declares that the function has no arguments (i.e., nargs was 0), then avalues is ignored. Note that argument values may be modified by the callee (for instance, structs passed by value); the burden of copying pass-by-value arguments is placed on the caller.

Note that while the return value must be register-sized, arguments should exactly match their declared type. For example, if an argument is a short, then the entry in avalues should point to an object declared as short; but if the return type is short, then rvalue should point to an object declared as a larger type – usually ffi_arg.

2.2 Simple Example

Here is a trivial example that calls puts a few times.

#include <stdio.h>
#include <ffi.h>
int main()
{
ffi_cif cif;
ffi_type *args[1];
void *values[1];
char *s;
ffi_arg rc;
/* Initialize the argument info vectors */
args[0] = &ffi_type_pointer;
values[0] = &s;
/* Initialize the cif */
if (ffi_prep_cif(&cif, FFI_DEFAULT_ABI, 1,
     &ffi_type_sint, args) == FFI_OK)
  {
    s = "Hello World!";
    ffi_call(&cif, puts, &rc, values);
    /* rc now holds the result of the call to puts */
    /* values holds a pointer to the function's arg, so to
       call puts() again all we need to do is change the
       value of s */
    s = "This is cool!";
    ffi_call(&cif, puts, &rc, values);
  }
return 0;
}

2.3 Types

2.3.1 Primitive Types

Libffi provides a number of built-in type descriptors that can be used to describe argument and return types:

ffi_type_void

The type void. This cannot be used for argument types, only for return values.

ffi_type_uint8

An unsigned, 8-bit integer type.

ffi_type_sint8

A signed, 8-bit integer type.

ffi_type_uint16

An unsigned, 16-bit integer type.

ffi_type_sint16

A signed, 16-bit integer type.

ffi_type_uint32

An unsigned, 32-bit integer type.

ffi_type_sint32

A signed, 32-bit integer type.

ffi_type_uint64

An unsigned, 64-bit integer type.

ffi_type_sint64

A signed, 64-bit integer type.

ffi_type_float

The C float type.

ffi_type_double

The C double type.

ffi_type_uchar
The C unsigned char type.
ffi_type_schar

The C signed char type. (Note that there is not an exact equivalent to the C char type in libffi; ordinarily you should either use ffi_type_schar or ffi_type_uchar depending on whether char is signed.)

ffi_type_ushort
The C unsigned short type.
ffi_type_sshort

The C short type.

ffi_type_uint
The C unsigned int type.
ffi_type_sint

The C int type.

ffi_type_ulong
The C unsigned long type.
ffi_type_slong

The C long type.

ffi_type_longdouble

On platforms that have a C long double type, this is defined. On other platforms, it is not.

ffi_type_pointer

A generic void * pointer. You should use this for all pointers, regardless of their real type.

ffi_type_complex_float
          The C _Complex float type.
ffi_type_complex_double
          The C _Complex double type.
ffi_type_complex_longdouble

The C _Complex long double type. On platforms that have a C long double type, this is defined. On other platforms, it is not.

Each of these is of type ffi_type, so you must take the address when passing to ffi_

prep_cif.

2.3.2 Structures

‘libffi’ is perfectly happy passing structures back and forth. You must first describe the structure to ‘libffi’ by creating a new ffi_type object for it.

[Data type] ffi_type

The ffi_type has the following members:

     size_t size

This is set by libffi; you should initialize it to zero.

     unsigned short alignment

This is set by libffi; you should initialize it to zero.

     unsigned short type

For a structure, this should be set to FFI_TYPE_STRUCT.

     ffi_type **elements

This is a ‘NULL’-terminated array of pointers to ffi_type objects. There is one element per field of the struct. Note that ‘libffi’ has no special support for bit-fields. You must manage these manually.

   The size and alignment fields will be filled in by ffi_prep_cif or ffi_prep_cif_var,

as needed.

2.3.3 Size and Alignment

libffi will set the size and alignment fields of an ffi_type object for you. It does so using its knowledge of the ABI.

You might expect that you can simply read these fields for a type that has been laid out by libffi. However, there are some caveats.

• The size or alignment of some of the built-in types may vary depending on the chosen ABI.

• The size and alignment of a new structure type will not be set by libffi until it has

been passed to ffi_prep_cif or ffi_get_struct_offsets.

• A structure type cannot be shared across ABIs. Instead each ABI needs its own copy of the structure type.

So, before examining these fields, it is safest to pass the ffi_type object to ffi_prep_ cif or ffi_get_struct_offsets first. This function will do all the needed setup.

ffi_type *desired_type;
ffi_abi desired_abi;
...
ffi_cif cif;
if (ffi_prep_cif (&cif, desired_abi, 0, desired_type, NULL) == FFI_OK)
  {
    size_t size = desired_type->size;
    unsigned short alignment = desired_type->alignment;
  }

libffi also provides a way to get the offsets of the members of a structure.

[Function] ffi_status ffi get struct offsets (ffi abi abi, ffi type *struct type, size t

*offsets) Compute the offset of each element of the given structure type. abi is the ABI to use; this is needed because in some cases the layout depends on the ABI. offsets is an out parameter. The caller is responsible for providing enough space for all the results to be written – one element per element type in struct type. If offsets is NULL, then the type will be laid out but not otherwise modified. This can be useful for accessing the type’s size or layout, as mentioned above. This function returns FFI_OK on success; FFI_BAD_ABI if abi is invalid; or FFI_BAD_ TYPEDEF if struct type is invalid in some way. Note that only FFI_STRUCT types are valid here.

2.3.4 Arrays, Unions, and Enumerations

2.3.4.1 Arrays

‘libffi’ does not have direct support for arrays or unions. However, they can be emulated using structures.

To emulate an array, simply create an ffi_type using FFI_TYPE_STRUCT with as many members as there are elements in the array.

ffi_type array_type;
ffi_type **elements
int i;
elements = malloc ((n + 1) * sizeof (ffi_type *));
for (i = 0; i < n; ++i)
  elements[i] = array_element_type;

elements[n] = NULL;
array_type.size = array_type.alignment = 0;
array_type.type = FFI_TYPE_STRUCT;
array_type.elements = elements;

Note that arrays cannot be passed or returned by value in C – structure types created like this should only be used to refer to members of real FFI_TYPE_STRUCT objects.

However, a phony array type like this will not cause any errors from ‘libffi’ if you use it as an argument or return type. This may be confusing.

2.3.4.2 Unions

A union can also be emulated using FFI_TYPE_STRUCT. In this case, however, you must make sure that the size and alignment match the real requirements of the union.

One simple way to do this is to ensue that each element type is laid out. Then, give the new structure type a single element; the size of the largest element; and the largest alignment seen as well.

This example uses the ffi_prep_cif trick to ensure that each element type is laid out.

ffi_abi desired_abi;
ffi_type union_type;
ffi_type **union_elements;
int i;
ffi_type element_types[2];
element_types[1] = NULL;
union_type.size = union_type.alignment = 0;
union_type.type = FFI_TYPE_STRUCT;
union_type.elements = element_types;
for (i = 0; union_elements[i]; ++i)
  {
    ffi_cif cif;
    if (ffi_prep_cif (&cif, desired_abi, 0, union_elements[i], NULL) == FFI_OK)
      {
        if (union_elements[i]->size > union_type.size)
          {
            union_type.size = union_elements[i];
            size = union_elements[i]->size;
          }
        if (union_elements[i]->alignment > union_type.alignment)
          union_type.alignment = union_elements[i]->alignment;
      }
  }

2.3.4.3 Enumerations

libffi does not have any special support for C enums. Although any given enum is implemented using a specific underlying integral type, exactly which type will be used cannot be determined by libffi – it may depend on the values in the enumeration or on compiler flags such as ‘-fshort-enums’. See Section “Structures unions enumerations and bit-fields implementation” in gcc, for more information about how GCC handles enumerations.

2.3.5 Type Example

The following example initializes a ffi_type object representing the tm struct from Linux’s

‘time.h’.

Here is how the struct is defined:

struct tm {
int tm_sec;
int tm_min;
int tm_hour;
int tm_mday;
int tm_mon;
int tm_year;
int tm_wday;
int tm_yday;
int tm_isdst;
/* Those are for future use. */
long int __tm_gmtoff__;
__const char *__tm_zone__;
};

Here is the corresponding code to describe this struct to libffi:

{
  ffi_type tm_type;
  ffi_type *tm_type_elements[12];
  int i;
  tm_type.size = tm_type.alignment = 0;
  tm_type.type = FFI_TYPE_STRUCT;
  tm_type.elements = &tm_type_elements;
  for (i = 0; i < 9; i++)
      tm_type_elements[i] = &ffi_type_sint;
  tm_type_elements[9] = &ffi_type_slong;
  tm_type_elements[10] = &ffi_type_pointer;
  tm_type_elements[11] = NULL;
  /* tm_type can now be used to represent tm argument types and
return types for ffi_prep_cif() */
}

2.3.6 Complex Types

‘libffi’ supports the complex types defined by the C99 standard (_Complex float, _

Complex double and _Complex long double with the built-in type descriptors ffi_type_
complex_float, ffi_type_complex_double and ffi_type_complex_longdouble.

Custom complex types like _Complex int can also be used. An ffi_type object has to be defined to describe the complex type to ‘libffi’.

[Data type] ffi_type

size_t size

This must be manually set to the size of the complex type.

unsigned short alignment

This must be manually set to the alignment of the complex type.

unsigned short type

For a complex type, this must be set to FFI_TYPE_COMPLEX.

ffi_type **elements

This is a ‘NULL’-terminated array of pointers to ffi_type objects. The first element is set to the ffi_type of the complex’s base type. The second element must be set to NULL.

The section Section 2.3.7 [Complex Type Example], page 9 shows a way to determine the size and alignment members in a platform independent way.

For platforms that have no complex support in libffi yet, the functions ffi_prep_cif and ffi_prep_args abort the program if they encounter a complex type.

2.3.7 Complex Type Example

This example demonstrates how to use complex types:

#include <stdio.h>
#include <ffi.h>
#include <complex.h>
void complex_fn(_Complex float cf,
                _Complex double cd,
                _Complex long double cld)
{
  printf("cf=%f+%fi\ncd=%f+%fi\ncld=%f+%fi\n",
         (float)creal (cf), (float)cimag (cf),
         (float)creal (cd), (float)cimag (cd),
         (float)creal (cld), (float)cimag (cld));
}
int main()
{
  ffi_cif cif;
  ffi_type *args[3];
  void *values[3];

_Complex float cf;
_Complex double cd;
_Complex long double cld;
/* Initialize the argument info vectors */
args[0] = &ffi_type_complex_float;
args[1] = &ffi_type_complex_double;
args[2] = &ffi_type_complex_longdouble;
values[0] = &cf;
values[1] = &cd;
values[2] = &cld;
/* Initialize the cif */
if (ffi_prep_cif(&cif, FFI_DEFAULT_ABI, 3,
                 &ffi_type_void, args) == FFI_OK)
  {
    cf = 1.0 + 20.0 * I;
    cd = 300.0 + 4000.0 * I;
    cld = 50000.0 + 600000.0 * I;
    /* Call the function */
    ffi_call(&cif, (void (*)(void))complex_fn, 0, values);
  }
return 0;
}

This is an example for defining a custom complex type descriptor for compilers that support them:

/*
* This macro can be used to define new complex type descriptors
* in a platform independent way.
*
* name: Name of the new descriptor is ffi_type_complex_<name>.
* type: The C base type of the complex type.
*/
#define FFI_COMPLEX_TYPEDEF(name, type, ffitype)             \
static ffi_type *ffi_elements_complex_##name [2] = {      \
  (ffi_type *)(&ffitype), NULL                             \
};                                                        \
struct struct_align_complex_##name {                      \
  char c;                                                  \
  _Complex type x;                                         \
};                                                        \
ffi_type ffi_type_complex_##name = {                      \
  sizeof(_Complex type),                                   \
  offsetof(struct struct_align_complex_##name, x),         \
  FFI_TYPE_COMPLEX,                                        \

         (ffi_type **)ffi_elements_complex_##name                 \
       }
     /* Define new complex type descriptors using the macro: */
     /* ffi_type_complex_sint */
     FFI_COMPLEX_TYPEDEF(sint, int, ffi_type_sint);
     /* ffi_type_complex_uchar */
     FFI_COMPLEX_TYPEDEF(uchar, unsigned char, ffi_type_uint8);

The new type descriptors can then be used like one of the built-in type descriptors in the previous example.

2.4 Multiple ABIs

A given platform may provide multiple different ABIs at once. For instance, the x86 platform has both ‘stdcall’ and ‘fastcall’ functions.

libffi provides some support for this. However, this is necessarily platform-specific.

2.5 The Closure API

libffi also provides a way to write a generic function – a function that can accept and decode any combination of arguments. This can be useful when writing an interpreter, or to provide wrappers for arbitrary functions.

This facility is called the closure API. Closures are not supported on all platforms; you can check the FFI_CLOSURES define to determine whether they are supported on the current platform.

Because closures work by assembling a tiny function at runtime, they require special allocation on platforms that have a non-executable heap. Memory management for closures is handled by a pair of functions:

[Function] void *ffi closure alloc (size t size, void **code)

Allocate a chunk of memory holding size bytes. This returns a pointer to the writable address, and sets *code to the corresponding executable address. size should be sufficient to hold a ffi_closure object.

[Function] void ffi closure free (void *writable)

Free memory allocated using ffi_closure_alloc. The argument is the writable address that was returned.

Once you have allocated the memory for a closure, you must construct a ffi_cif describing the function call. Finally you can prepare the closure function:

[Function] ffi_status ffi prep closure loc (ffi closure *closure, ffi cif *cif, void (*fun) (ffi cif *cif, void *ret, void **args, void *user_data), void

        *user_data, void *codeloc)

Prepare a closure function. The arguments to ffi_prep_closure_loc are:

closure The address of a ffi_closure object; this is the writable address re-

               turned by ffi_closure_alloc.

cif The ffi_cif describing the function parameters. Note that this object, and the types to which it refers, must be kept alive until the closure itself is freed.

user data An arbitrary datum that is passed, uninterpreted, to your closure function.

codeloc The executable address returned by ffi_closure_alloc.

fun The function which will be called when the closure is invoked. It is called with the arguments:

          cif  The ffi_cif passed to ffi_prep_closure_loc.

ret A pointer to the memory used for the function’s return value. If the function is declared as returning void, then this value is garbage and should not be used. Otherwise, fun must fill the object to which this points, following the same special promotion behavior as ffi_call. That is, in most cases, ret points to an object of exactly the size of the type specified when cif was constructed. However, integral types narrower than the system register size are widened. In these cases your program may assume that ret points to an ffi_arg object.

args A vector of pointers to memory holding the arguments to the function.

user data The same user data that was passed to ffi_prep_closure_

                    loc.

ffi_prep_closure_loc will return FFI_OK if everything went ok, and one of the other ffi_status values on error. After calling ffi_prep_closure_loc, you can cast codeloc to the appropriate pointer-to-function type.

You may see old code referring to ffi_prep_closure. This function is deprecated, as it cannot handle the need for separate writable and executable addresses.

2.6 Closure Example

A trivial example that creates a new puts by binding fputs with stdout.

#include <stdio.h>
#include <ffi.h>
/* Acts like puts with the file given at time of enclosure. */
void puts_binding(ffi_cif *cif, void *ret, void* args[],
                  void *stream)
{
  *(ffi_arg *)ret = fputs(*(char **)args[0], (FILE *)stream);
}

typedef int (*puts_t)(char *);
int main()
{
ffi_cif cif;
ffi_type *args[1];
ffi_closure *closure;
void *bound_puts;
int rc;
/* Allocate closure and bound_puts */
closure = ffi_closure_alloc(sizeof(ffi_closure), &bound_puts);
if (closure)
  {
    /* Initialize the argument info vectors */
    args[0] = &ffi_type_pointer;
    /* Initialize the cif */
    if (ffi_prep_cif(&cif, FFI_DEFAULT_ABI, 1,
                     &ffi_type_sint, args) == FFI_OK)
      {
        /* Initialize the closure, setting stream to stdout */
        if (ffi_prep_closure_loc(closure, &cif, puts_binding,
                                 stdout, bound_puts) == FFI_OK)
          {
            rc = ((puts_t)bound_puts)("Hello World!");
            /* rc now holds the result of the call to fputs */
          }
      }
  }
/* Deallocate both closure, and bound_puts */
ffi_closure_free(closure);
return 0;
}

2.7 Thread Safety

libffi is not completely thread-safe. However, many parts are, and if you follow some simple rules, you can use it safely in a multi-threaded program.

• ffi_prep_cif may modify the ffi_type objects passed to it. It is best to ensure that only a single thread prepares a given ffi_cif at a time.

• On some platforms, ffi_prep_cif may modify the size and alignment of some types, depending on the chosen ABI. On these platforms, if you switch between ABIs, you must ensure that there is only one call to ffi_prep_cif at a time. Currently the only affected platform is PowerPC and the only affected type is long

  double.

3 Memory Usage

Note that memory allocated by ffi_closure_alloc and freed by ffi_closure_free does not come from the same general pool of memory that malloc and free use. To accomodate security settings, ‘libffi’ may aquire memory, for example, by mapping temporary files into multiple places in the address space (once to write out the closure, a second to execute it). The search follows this list, using the first that works:

• A anonymous mapping (i.e. not file-backed)

• memfd_create(), if the kernel supports it.

• A file created in the directory referenced by the environment variable LIBFFI_TMPDIR.

• Likewise for the environment variable TMPDIR.

• A file created in /tmp.

• A file created in /var/tmp.

• A file created in /dev/shm.

• A file created in the user’s home directory ($HOME).

• A file created in any directory listed in /etc/mtab.

• A file created in any directory listed in /proc/mounts.

If security settings prohibit using any of these for closures, ffi_closure_alloc will fail.

4 Missing Features

libffi is missing a few features. We welcome patches to add support for these. • Variadic closures.

• There is no support for bit fields in structures.

• The “raw” API is undocumented.

• The Go API is undocumented.

Index 15

Index

A

ABI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Application Binary Interface . . . . . . . . . . . . . . . . . . . . 1

C

calling convention. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 cif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 closure API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 closures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

F

FFI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ffi_call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ffi_closure_alloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ffi_closure_free. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 FFI_CLOSURES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ffi_get_struct_offsets . . . . . . . . . . . . . . . . . . . . . . . 6 ffi_prep_cif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ffi_prep_cif_var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ffi_prep_closure_loc . . . . . . . . . . . . . . . . . . . . . . . . 11 ffi_status . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 2, 6, 11 ffi_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 9

ffi_type_complex_double . . . . . . . . . . . . . . . . . . . . . . 5

ffi_type_complex_float . . . . . . . . . . . . . . . . . . . . . . . 5

ffi_type_complex_longdouble . . . . . . . . . . . . . . . . . 5

ffi_type_double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_longdouble . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ffi_type_pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ffi_type_schar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_sint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_sint16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_sint32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_sint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_sint8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_slong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_sshort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_uchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_uint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_uint16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_uint32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_uint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_uint8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_ulong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_ushort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ffi_type_void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Foreign Function Interface . . . . . . . . . . . . . . . . . . . . . . 1

V

void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 11